Point charge q_1 = +3.67 mu C is at point A (-37.2
Answer #1

Field due to +3.67 uC is E1 = k*q/x^2 = (9*10^9*3.67*10^-6)/(0.372^2) = 2.38*10^5 N/C

Field due to -3.67 uC is E2 = k*q/x^2 = (9*10^9*3.67*10^-6)/(0.372^2) = 2.38*10^5 N/C


these E1 and E2 are in same direction,hence Enet = E1+E2 = 2*2.38*10^5 = 4.76*10^5 N/C

so a) E = 4.76*10^5 N/C


b) toward the negative charge

c) net force on electron is Fnet = q*E = 1.6*10^-19*4.76*10^5

Fnet = 7.616*10^-14 N

but Fnet = m*a

m*a = 7.616*10^-14

accelaration a = 7.616*10^-14/(9.11*10^-31) = 8.36*10^16 m/s^2

the direction of force on electron is always opposite to E

hence the direction of force is toward the positive charge